Prev Next

SW&IT Meeting Minutes 2020-11-16

Attendees: Carl, David, Igor, Matthieu, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20201102
LR20201102 - Bec, Matthieu 72.0
LR20201102 - Holmberg, Carl 72.0
LR20201102 - Edmeades, David 72.0
LR20201102 - Sola, Igor 72.0
LR20201102 - Grenz, Paul 20.0
LR20201102 - Hooper, Stephen 80.0
LR20201102 - Bryant, Trey 0.0
LR20201102 - Zhang, Yang 72.0
LR20201102 - Zhang, Xianyu 72.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT
    • DMS AO msql adapter
    • cobbler pxe/kickstart
  • This week
    • ATT SW, ATT PM monthly, ROWG, SHARK-VIS waffle, AO MCS SW meetings
    • DMS AO discussion with SOUL team

DE

  • Synology says that we need to update DSM on mountain unit; on hold until safe time and LD up
  • Office visit
    • Couple of desktop physical issues solved
    • Server room UPS plan
  • Cisco
This week:
  • Thursday meetings
    • IBM storage
    • UITS switch
  • DT IRTC computer multiboot
  • UPS battery purchase?
  • News
    • Not recommending upgrading OSX; reports of bricking older models and software incompatibilities
    • Win 10 password issue with MS workaround
  • Cisco
  • ATTWG Zoom recordings: No one has responded to my offer to forward the links anywhere specific, but I have archived them. If anyone is wondering how to access the recordings let me know.
  • Archive machine resource report

CH

  • Last week:
    • AGW:
      • Pushed new PMAC 'firmware' from JP into github. Is it AGW1-specific? If not, look at renaming.
      • Contacted Delta Tau to revisit (2009 v. 2020) the possibility of adding a 'pmac' firmware upload capability.
        • Reply: PMAC command/comms protocol (still) doesn't support a firmware download capability outside of PEWIN.
      • Tested encoder with manual turns to check the value is correct.
        • Compared register values on mock UMAC and actual AGW1 using the firmware formula to calculate M162 from M101. AGW1 is good, mock UMAC not.
        • Recorded register data from 'pmac' client during manual 90 and 180 degree-ish turns with the binary encoder, forwarded results to JP.
        • Recorded register data from PEWIN32, forwarded results to JP. His analysis is that the encoder/serial interface are returning valid results.
        • JP's analysis is that the mock UMAC firmware is not dropping into the if/then branches surrounding the desired M101 to M162 calculation. Manually set the controlling registers (Y_HOMEOFFSET, MOTION_IN_PROGRESS), but no change. JP will continue to evaulate the code vs. results.
    • AHS: Fixed byte stream to string conversion error, caused by change in LBTI INDI parameter.
      • (temp_ctrl2.sensor.enable_collection replaces temp_ctrl2.sensor.notify_out_of_range)
  • This week:
    • AGW:
      • Follow up with JP on encoder calculations within firmware.
      • Follow up with Delta Tau to revisit 'pmac' upload capability.
    • AHS: Add collection of _STATE parameter to validate reading Cryomech ADAM data from LBTI INDI service. Add initialization parameter status check.
    • statserv: investigate and test which INDI drivers, if any are explicitly required, to restore web page updates.

YZ

  • Previous:
    • TMS test and data analysis, meeting etc.
    • Some TCS work (AOS and DDSGUI patches).
    • Some AO-Cravat/OSCO work.
    • Beginning modsAlign refactoring.
    • Completed UA harassment training.
  • This week:
    • Usual TMS and AO-Cravat/OSCO work.
    • Continue MODS programs refactoring.

XZ

  • Last week
    • AOS changes for ASM elevation limit display
    • SX AdSec retracted test
    • ASM DSP firmware refresh lab setup plan/meeting
    • AOS/ALH for TSS display invistigation
    • ATT AdSec
    • ATT SW
    • AO WFS mysql
    • AO operation support
    • few sections of UA harassment trainig
  • This week
    • ATT AdSec
    • ATT SW
    • Completed UA harassment training
    • AOS ASM elevation limit display changes on AO supervisor side
    • AO operation support
    • AOS/ALH for TSS display modification

IS

Last week:
  • LUCI
    • RPMs generation. Two RPMs configuration and byte code
    • Geirs script. New way of observing, not done before
  • PIT
    • pit admin tool

This week:
  • Harassment training
  • LUCI
    • RPM
    • IT 8231 GEIRS timeout
    • IT 8220
    • split current luci motor privilege into two to allow more control on what people can and can't do.
    • Look into reading WebIO status and passing it to the database/telemetry.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
  • PIT
    • pit admin tool

PSG

Last Week:
  • Researched existing code for ARGOS Swing arm and generated design document.
This Week:
  • Start to code based on design.

SH

  • HTTPS info
  • obs5 and obs6 installation of MATE
  • Update VPN clients
  • info redirects
  • Joe and JP's ethernet
  • Zabbix work

next:

  • schedule measurement of mountain bandwidth
  • Wired registration switch meeting (UITS)
  • ROWG (monitors)
  • PEPSI with Leroy
  • Scripting for SSL Let's Encrypt
Topic revision: r14 - 16 Nov 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback